# QAEncoder: Towards Aligned Representation Learning in Question Answering System

### v1

@@ 4.1.2. Experiments On Latest Datasets
As illustrated in Table 2, for the latest datasets, QAEncoder significantly enhances the alignment of user queries with relevant documents across both state-of-the-art embedding models and other popular open-source models. For instance, the gte-multilingual-base model’s MRR metric on the FIGNEWS(English) dataset increases from 65.3 to 75.3. Similarly, the mContriever model’s MRR on the FIGNEWS(English) dataset improves from 32.8 to 61.2. Besides, the text2vec-base-multilingual model’s MRR on the CRUD-RAG dataset rose from 12.1 to 55.3. These results confirm QAEncoder’s remarkably generalized alignment capability across various embedding models and multilingual datasets.
4.2. Analysis and Discussion
@@
For a more comprehensive assessment, we also report results of various QAEncoder ablations, i.e. QAEemb, QAEtxt, QAEhyb and QAEnaive which directly stores predicted queries. Finally, we also discuss the relationship between QAEncoder and training-based as well as document-centric methods.
4.2.1. Ablation Studies
@@
Table 3. Performance comparison of QAEncoder variants on the latest datasets FIGNEWS and CRUD-RAG. Higher is better, the best one is bolded. Hyperparameters are optimized simultaneously across all the latest datasets. $n$ indicates the number of predicted queries in QAnaive.
Model Param FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese)
MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8
bge-m3 QAEemb, $\alpha=0.3$ 76.2 80.0 80.0 82.6 88.7 90.6
QAEtxt, $\beta=1.5$ 77.1 80.8 80.1 82.8 89.0 90.8
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 77.3 80.7 80.6 83.3 89.4 91.2
QAEnaive, $n=10$ 76.8 79.5 76.9 79.2 85.9 87.8
multilingual-e5-large QAEemb, $\alpha=0.45$ 77.8 81.1 79.7 82.7 89.7 91.4
QAEtxt, $\beta=1.5$ 75.6 79.0 80.7 83.7 88.3 90.4
QAEhyb, $\alpha=0.15,\;\beta=1.25$ 77.0 80.5 82.1 84.9 89.1 91.1
QAEnaive, $n=10$ 76.8 79.4 76.5 79.2 84.9 86.8
@@
We present the performance comparison of QAEncoder variants on two state-of-the-art multilingual embedding models in Table 3. Generally, QAEhyb and QAEemb outperform the QAEtxt and QAEnaive approaches. For instance, for the bge-m3 model, QAEhyb consistently outperforms other variants. Conversely, the multilingual-e5-large model performs best with QAEemb. However, the best performance differences between QAEemb, QAEtxt, and QAEhyb are not substantial, demonstrating the robustness of our approach to hyperparameter variations. Regarding QAEnaive, it evidently underperforms other ablations, despite storing 10 times the number of embedding vectors. This leads to unacceptable storage management overhead and recall latency in large-scale production systems. We also provide some more granular ablation experiments in Figure 4 and the appendix.
Figure 4. Ablation on $\alpha$ and $n$ hyperparameters for QAEemb and QAEnaive on FIGNEWS(English) dataset.
4.2.2. Training-based and Document-centric Methods
Training-based approaches mainly include two categories, fine-tuning on QA datasets and multi-task instruction datasets. We choose mContriever-msmarco and multilingual-e5-large-instruct as representative models respectively. Finally, we consider HyDE (Gao et al., 2023), a well-known document-centric method for comparison.
Table 4. Performance comparison of QAEncoder against training-based and document-centric methods on the latest datasets FIGNEWS and CRUD-RAG. Higher is better, the best one is bolded. Hyperparameters $\alpha,\;\beta$ are optimized simultaneously across the two latest datasets. $n$ denotes the number of pseudo-documents in HyDE. - indicates default or null values.
Model Param FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese)
MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8
mContriever-msmarco - 65.9 70.0 70.1 73.5 85.0 87.4
QAEhyb, $\alpha=0.3,\;\beta=0.75$ 72.1 76.3 77.2 80.2 88.6 90.6
multilingual-e5-large-instruct - 66.7 70.7 74.9 78.1 79.8 82.5
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 75.5 79.4 80.7 83.5 88.9 90.6
mContriever - 32.8 36.3 40.1 44.2 71.7 75.8
HyDE, $n=8$ 24.9 27.6 35.1 40.4 70.7 74.2
multilingual-e5-large - 73.8 77.5 76.6 79.9 85.7 88.2
HyDE, $n=8$ 63.4 67.7 68.2 73.8 81.5 83.9
@@ 4.2.2. Training-based and Document-centric Methods
Training-based and query-centric methods operate at training time and indexing time, respectively. Therefore, integrating these approaches cloud lead to more improvements. As illustrated in Table 4, both types of fine-tuned models significantly benefit from the QAEncoder. For instance, the mContriever-msmarco model improves MRR from 70.1 to 77.2 on FIGNEWS(Arabic); the multilingual-e5-large-instruct model’s MRR increases 8.8 and 9.1 MRR points on the FIGNEWS(English) and CRUD-RAG(Chinese) datasets, respectively.
@@
For document-centric methods, we investigate the widely-reported hallucination phenomenon on latest datasets (Wang et al., 2023a; Gao et al., 2023; Kim and Min, 2024). Table 4 shows the recall performance heavily decreases for all the latest datasets and both models, attributed to the hallucination of pseudo-documents. Besides, the LLM invocation for pseudo-documents is both time-consuming and costly. In our case, the time for single LLM invocation is greater than 2000ms while the time for vector search is less than 10ms. These highlight the irreplaceable importance and practicality of QAEncoder method.
5. Conclusion
In this paper, we introduced QAEncoder, a training-free approach to bridge the document-query gap for more advanced QA systems. Based on our conical distribution hypothesis and document fingerprint strategies, QAEncoder substitutes document embeddings with the expectation of query embeddings, enriched with document information. Extensive experiments across multiple datasets, languages and embedding models confirm that QAEncoder not only avoids extra index storage, retrieval latency, training costs, and hallucination, but also ensures robust generalization and diversified query handling.
Appendix A Appendix
A.1. Dataset Details
  • MS-MARCO is a large-scale dataset specifically designed for machine reading comprehension. It comprises real user queries collected through Bing search, paired with corresponding passages retrieved from a comprehensive web document collection. We evaluated our model on the passage retrieval task using a corpus of approximately 8.8M passages.
  • Natural Questions (NQ) is a widely used dataset comprising real user queries submitted to the Google search engine, paired with relevant Wikipedia pages. Each query is annotated with a long answer (typically a paragraph) and a short answer (one or more entities). Each question-answer pair corresponds to a single Wikipedia page, ensuring clarity in evaluating information retrieval performance.
  • TriviaQA is a dataset of question-answer-evidence triples, with over 650K entries, including 95K question-answer pairs created by trivia enthusiasts. The evidence documents, typically sourced from Wikipedia or other web sources, average six per question.
  • HotPotQA contains 113K Wikipedia-based question-answer pairs that require multi-hop reasoning across multiple pages. Each pair includes a set of supporting sentences, which are treated as evidence documents.
  • ELI5 is a question-answering dataset consisting of question-answer-evidence triples, where the questions are complex and often require detailed, multi-sentence responses. It contains 270K diverse and intricate questions that necessitate explanatory answers. To support each response, web search results are utilized as evidence documents, enhancing the reliability of the provided answers.
  • SQuAD is a widely used reading comprehension dataset comprising 107K question-answer pairs derived from 536 Wikipedia articles. Each question are specific text segments from the relevant articles. In this dataset, the Wikipedia articles function as evidence documents, supplying the essential context required for accurate response retrieval.
  • CRUD-RAG is a benchmark specifically designed for evaluating RAG systems, incorporating the latest high-quality news data that were not included in the training phase of the language models. It comprises more than 80K news articles sourced from prominent Chinese news websites, all published after July 2023. From the set of queries generated by GPT-4o-mini for each document, we randomly sample one to serve as the test query. The original news is designated as the evidence documents for recall evaluation, ensuring queries are associated with exactly relevant documents. [^1] LlamaIndex adopts the same common practice and provides templated workflow, i.e. generating queries for recall and rerank evaluation. See https://www.llamaindex.ai/blog/boosting-rag-picking-the-best-embedding-reranker-models-42d079022e83.
A.2. Metric Details
For SQuAD, CRUD-RAG and FIGNEWS benchmarks, MAP metric is equivalent to MRR, as there is only one related document per query.
  • Mean Reciprocal Rank (MRR) Mean Reciprocal Rank is a statistic measure used to evaluate the effectiveness of a retrieval system by calculating the reciprocal of the rank at which the first relevant result appears. The mathematical formulation is:
    $\text{MRR}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\frac{1}{\text{rank}_{i}}$
  • Mean Average Precision (MAP) Mean Average Precision is a measure that combines precision and recall. It computes the average precision value for a set of queries and then averages these values. The mathematical formulation is:
    $\text{MAP}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\text{AP}(i),$
    Where: - $|Q|$ is the number of queries. - $\text{AP}(i)$ is the average precision for the $i$-th query.
    The average precision for a single query is given by:
    $\text{AP}=\frac{1}{R}\sum_{k=1}^{n}P(k)\cdot\text{rel}(k),$
    Where: - $R$ is the total number of relevant documents for the query. - $n$ is the number of retrieved documents. - $P(k)$ is the precision at cut-off $k$. - $\text{rel}(k)$ is an indicator function equating to 1 if the document at rank $k$ is relevant, otherwise 0.
  • Normalized Discounted Cumulative Gain (NDCG) Normalized Discounted Cumulative Gain is a measure of ranking quality that takes into account the positions of the relevant documents. It is based on the concept of discounting the relevance of documents based on their position in the result list. The mathematical formulation is:
    $\text{NDCG}=\frac{DCG_{p}}{IDCG_{p}},$
    The Discounted Cumulative Gain at position $p$ is given by:
    $DCG_{p}=\sum_{i=1}^{p}\frac{2^{\text{rel}_{i}}-1}{\log_{2}(i+1)},$
    Where: - $\text{rel}_{i}$ is the relevance score of the document at rank $i$.
A.3. Baseline Details
  • DeepImpact (Mallia et al., 2021b) leverages DocT5Query to enrich the document collection and estimates the semantic importance of tokens with fine-tuned BERT.
  • BGE models (Xiao et al., 2024)by BAAI, Jina models (Günther et al., 2023) by Jina AI, E5 models (Wang et al., 2023b) by Microsoft, GTE models (Zhang et al., 2024b) by Alibaba-NLP are the most advanced embedding models, featuring multilingual understanding and task-specific instruction tuning capabilities. We choose both vanilla encoders and QA-specific instruction-tuned encoders for test.
  • Training-based approaches mainly include two categories, fine-tuning on QA datasets or multi-task instruction datasets. For fine-tuning on QA datasets, we choose mContriever-msmarco for test, an enhanced variant of the mContriever model that has been fine-tuned on the MS-MARCO. The second category involves fine-tuning models on multi-task instruction datasets, where distinct prompt prefixes are appended to the input text, enabling the model to effectively differentiate between various tasks. In this category, we test the multilingual-e5-large-instruct (Wang et al., 2024b) developed by Microsoft, which leverages synthetic instruction data (Wang et al., 2024a) for fine-tuning.
  • Document-centric methods instruct LLMs to generate a pseudo-document for each query. The pseudo-document aims to capture relevant information but does not correspond to a real document and may contain inaccuracies and hallucinations. Subsequently,the pseudo-document is encoded, and its embedding is utilized to retrieve similar real documents based on vector similarity. We choose Hyde (Gao et al., 2023) for test.
A.4. Prompts
We employ specialized prompts to instruct gpt-4o-mini as the question and pseudo-document generator. Only the English version is presented due to LaTeX compilation issues with non-English languages.
A.5. Proof of the Conical Distribution Hypothesis
@@
A.6. Strong Conical Distribution Hypothesis
In this subsection, we further substantiate the original hypothesis that the potential queries adhere to a Gaussian distribution: For any document $d$, the potential queries in the embedding space approximately follow a Gaussian distribution, characterized by a mean $\mu$ and covariance matrix $\Sigma$. Refer to A.6.3 for detailed validation.
A.6.1. Main theorem

### v3

@@ 4.1.2 Performance on Latest Datasets
In scenarios such as search engine, financial analysis, and news QA, large volumes of new data constantly emerge and are indexed into retrieval base for accurate and up-to-date response. Hence, the alignment capability for previously unseen user queries and relevant documents is crucial for embedding models in RAG systems. We experiment on the latest news datasets, FIGNEWS and CRUD-RAG, to avoid data leakage and mimic the real-world scenarios. As illustrated in Table 2, for latest datasets, QAEncoder significantly improves across both state-of-the-art embedding models and other well-known models. E.g., the gte-multilingual-base model and the mcontriever model’s MRR metrics increase from 65.5 to 75.5, and 32.9 to 61.4 respectively on FIGNEWS(English) dataset. The multilingual-e5-large model’s MRR increase from 76.7 to 82.2 on FIGNEWS(Arabic). Additionally, the text2vec-base-multilingual model’s MRR on CRUD-RAG dataset rises from 9.7 to 32.1, while the bge-m3 model’s MRR improves from 47.5 to 51.4. These results remarkably confirm QAEncoder’s generalized alignment capability across various embedding models and languages. See Tables 8 and 11 for more embedding models and datasets.
4.2 Analysis and Discussion
@@
For more comprehensive assessments, we analyze various QAEncoder ablations, i.e. QAEemb, QAEtxt, QAEhyb and QAEnaive, which directly stores predicted queries. We also evaluate QAEncoder’s robustness with respect to the query generator. Finally, we discuss the relationship between QAEncoder and both training-based and document-centric methods.
4.2.1 Ablations of QAEncoder
Model Method FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese)
MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10
bge-m3 QAEemb, $\alpha=0.3$ 76.4 80.5 80.1 82.9 51.3 52.4
QAEtxt, $\beta=1.5$ 77.2 81 80.2 83.1 51.4 52.5
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 77.4 81.1 80.6 83.4 51.7 52.7
QAEnaive, n=10 76.9 79.9 77.1 79.7 47.0 48.1
multilingual-e5-large QAEemb, $\alpha=0.45$ 77.9 81.4 79.8 83 51.9 52.9
QAEtxt, $\beta=1.5$ 75.6 79.2 80.9 84.1 51.0 52.3
QAEhyb, $\alpha=0.15,\;\beta=1.25$ 77.1 80.6 82.2 85.1 51.5 52.7
QAEnaive, n=10 77.5 80.3 76.5 79.4 46.5 47.7
@@ 4.2.1 Ablations of QAEncoder
Table 3: Performance comparison of QAEncoder variants on the latest datasets. Hyperparameters are optimized simultaneously for all the latest datasets. $n$ indicates the number of predicted queries in QAEnaive.
@@
We present the performance comparison of QAEncoder variants on two state-of-the-art embedding models in Table 3. Generally, QAEhyb and QAEemb outperform the QAEtxt and QAEnaive approaches. For instance, for the bge-m3 model, QAEhyb consistently outperforms other variants. Conversely, the multilingual-e5-large model performs best with QAEemb. However, the best performance differences between QAEemb, QAEtxt, and QAEhyb are not substantial, demonstrating the robustness of our approach to hyperparameter variations. Regarding QAEnaive, it evidently underperforms other ablations, despite storing 10 times the number of embedding vectors. This leads to unacceptable storage management overhead and recall latency in large-scale production systems. We provide more granular ablation experiments in Fig. 4, as well as the convergence speed of QAEbase’s Monte Carlo estimation. See Table 9 for the full table.
Figure 4: Left: The impact of varying $\alpha$ values for QAEemb. Middle: The effect of varying the number of predicted queries for QAEnaive, with MRR improving as $n$ increases, approaching stability at $n=10$. The curves for different models are mostly identical as the matching are largely driven by predicted queries. Right: The convergence speed for Monte Carlo estimation of QAEbase. Please refer to Appendix C and Fig. 11 for more cost details.
4.2.2 Robustness w.r.t. Query Generator
[…]
As shown in Tab. 4, we adopted GPT-4o-mini to generate test queries and various LLMs as generators for query prediction. For QAEemb, GPT-4o-mini shows a slight advantage over other predictors. But for QAEtxt and QAEhyb, the performance is largely consistent, highlighting not only the importance of document fingerprints but also the robustness of QAEncoder. Interestingly, Claude beats GPT on QAEtxt with MRR values 73.5 and 72.7 respectively, while the vanilla BGE is 66.1. The full names of abbreviations are in left Appendix B.
Model Method GPT Claude Qwen DeepSeek
bge-large-en-v1.5 QAEemb, $\alpha=0.5$ 72.2 71.6 71.0 70.5
QAEtxt, $\beta=1.5$ 72.7 73.5 72.6 72.5
QAEhyb, $\alpha=0.15,\beta=1.5$ 73.7 73.5 73.2 73.3
multilingual-e5-large QAEemb, $\alpha=0.5$ 77.6 76.1 75.8 75.9
QAEtxt, $\beta=1.5$ 75.1 74.9 74.7 74.6
QAEhyb, $\alpha=0.15,\beta=1.5$ 77.0 76.5 76.6 76.6
Table 4: MRR@10 on FIGNEWS(English) with GPT-generated test queries and different query predictors. The performance is largely robust to query predictors.
Model Method FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese)
MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10
mcontriever - 32.9 36.7 40.3 44.7 39.2 41.6
QAEhyb, $\alpha=0.45,\;\beta=1.25$ 61.4 65.9 68.3 72.1 51.3 52.4
GPL 68.3 73.6 72.1 75.18 52.8 55.9
MS 66.1 70.6 70.2 73.7 46.5 47.8
MS + QAEhyb, $\alpha=0.3,\;\beta=0.75$ 72.3 76.8 77.3 80.5 51.2 52.4
QA-RAG 31.1 34.4 42.3 46.8 43.8 45.8
Query2Doc 25.1 29.0 34.5 38.9 35.7 37.2
HyDE 25 27.9 35.7 41.9 36.7 38.7
multilingual-e5-large - 73.9 77.8 76.7 80.2 46.9 48.3
QAEhyb, $\alpha=0.15,\;\beta=1.25$ 77.1 80.6 82.2 85.1 51.5 52.7
GPL 75.2 78.9 79.4 82.3 53.6 56.3
INS 67 71.4 75 78.2 43.7 45.2
INS + QAEhyb, $\alpha=0.15,\;\beta=1.5$ 75.6 79.8 80.8 83.7 51.4 52.4
QA-RAG 73.3 76.5 72.8 75.6 45.8 46.5
Query2Doc 63.4 68.2 66.3 72.8 42.0 43.1
HyDE 63.6 68.3 68.3 74.1 42.3 43.6
Table 5: Performance comparison of QAEncoder with training-based and document-centric methods on the latest datasets FIGNEWS and CRUD-RAG. Hyperparameters are optimized simultaneously for all latest datasets. $\dagger$ indicates the training-based methods. MS represents fine-tuning on MSMARCO, i.e. the mcontriever-msmarco model; INS represents instruction-tuning, i.e. the multilingual-e5-large-instruct model (Wang et al., 2023a). $\ddagger$ indicates the document-centric methods. See Table 10 for the full table.
4.2.3 Training-based and Document-centric Methods
Training-based approaches mainly include two types: fine-tuning on QA datasets (domain adaptation) and fine-tuning on multi-task instruction datasets. We select mcontriever-msmarco and multilingual-e5-large-instruct as representative methods respectively. To unveil the catastrophic forgetting issue of training-based methods, we incorporate GPL Wang et al. (2021), which predicts queries, mines hard negative samples, and distills the re-ranker for unsupervised domain adaptation.
For multi-domain adaptation, we fine-tune on FIGNEWS and CRUD-RAG datasets iteratively to simulate knowledge update. While GPL bridges the document-query gap with notable improvements on CRUD-RAG dataset, the catastrophic forgetting issue is serious: GPL improves marginally on FIGNEWS datasets, while QAEncoder works robustly due to the training-free nature.
@@ 4.2.3 Training-based and Document-centric Methods
Training-based and query-centric methods operate at training time and indexing time, respectively. Therefore, integrating these approaches could lead to more improvements. As illustrated in Table 5, both types of fine-tuned models significantly benefit from the QAEncoder. For instance, the mcontriever-msmarco model improves MRR from 70.2 to 77.3 on FIGNEWS(Arabic); the multilingual-e5-large-instruct model’s MRR increases 8.6 and 7.7 MRR points on the FIGNEWS(English) and CRUD-RAG(Chinese) datasets, respectively.
@@
For document-centric methods, we consider HyDE (Gao et al., 2023), Query2Doc (Wang et al., 2023b), and QA-RAG (Kim and Min, 2024) for comparison. Note that the pseudo-document generator in QA-RAG is fine-tuned on medical QA datasets, here we adopt out-of-box GPT-4o-mini. The widely-reported hallucination phenomenon on the latest datasets is confirmed (Wang et al., 2023b). As shown in Tab. 5, the retrieval performance of HyDE and Query2Doc heavily decreases for all the latest datasets, attributed to the hallucination of pseudo-document generation. Although QA-RAG alleviates hallucination by its two-way retrieval and re-ranking mechanism, it still underperforms. Besides, the LLM invocation for pseudo-documents is both costly and time-consuming. In our case, the time for single LLM invocation is more than 2000ms while the time for vector search is less than 10ms. These highlight the irreplaceable importance and practicality of QAEncoder method.
5 Conclusion
In this paper, we propose QAEncoder to bridge the document-query gap from the query-centric perspective—a novel, training-free and pioneering approach. QAEncoder replaces document embeddings with the expectation of query embeddings, theoretically supported by the conical distribution hypothesis and practically enhanced by document fingerprint strategies. Extensive experiments are conducted on both classical BEIR benchmark suite and the latest news datasets, covering 20+ embedding models and 6 languages. QAEncoder demonstrates robust generalization, diverse query handling, and compatibility with existing RAG architectures and training-based methods.
Acknowledgments
This work is supported by the National Key R&D Program of China (2024YFA1014003), National Natural Science Foundation of China (92470121, 62402016), CAAI-Ant Group Research Fund, and High-performance Computing Platform of Peking University.
Limitations
Despite the benefits of QAEncoder, there are still ongoing works for further improvements:
  • The current mean pooling could be overly simplistic and limits the performance improvement. The multi-cluster version such as Gaussian mixture models and multi-vector representation could be explored.
  • Since the query generator mainly generates simple queries, out-of-domain issues with complex, multi-hop queries could happen. However, query optimization has become increasingly prevalent to refine the original user query, making it more suitable for retrieval module. Practically, query decomposition and rewriting can be integrated to break down complex queries into simpler sub-queries and to rewrite these sub-queries into the in-domain style, respectively. Therefore, we focus QAEncoder on the core retrieval module, and leave the handling of complex queries to auxiliary strategies.
  • The 5W1H framework captures our core design principles and serves as the motivation for our work, primarily applicable to narrative or factual texts. Without loss of generality, QAEncoder estimates the cluster center of potential queries via Monte Carlo method, where the type of query is arbitrary and not constrained to 5W1H. We acknowledge there could be the worst case, user query may deviate from predicted ones, i.e. the outlier. Therefore, more fallback mechanisms beyond document fingerprints possess research value.
  • There is also a risk of data leakage for property information when predicting queries via API calls. The query prediction process may incur underlying security vulnerabilities.
We aim to keep QAEncoder simple-yet-effective, and leave these problems in the future research.
Appendix
[…]
Appendix A Conical Distribution Hypothesis
A.1 Case Studies
Here we provide several concrete case studies in Fig. 6, 7, 8, 9. The conclusions are highly consistent with experiments in the main text, across embedding models and languages. Please refer to our repository for related scripts and reproduction.
Figure 6: Similarity matrix and angle distribution on Armstrong example with contriever encoder.
Figure 7: Similarity matrix and angle distribution on Einstein example with contriever encoder.
Figure 8: Similarity matrix and angle distribution on Curie example with contriever encoder.
Figure 9: Similarity matrix and angle distribution on Armstrong example in French with multilingual-e5-large.
A.2 Proof of Conical Distribution Hypothesis
@@
Furthermore, when the stronger hypothesis assuming the cluster follows the Gaussian distribution is adopted, more quantitative analysis results can be derived.
A.3 Strong Conical Distribution Hypothesis
In this subsection, we further substantiate the original hypothesis that the potential queries adhere to a Gaussian distribution: For any document $d$, the potential queries in the embedding space approximately follow a Gaussian distribution, characterized by a mean $\mu$ and covariance matrix $\Sigma$. Refer to Appendix A.3.3 for detailed validation.
A.3.1 Main theorem